# Learning to Generate Text in Arbitrary Writing Styles

### 28 Dec 2023 v1

@@ 3.2 An energy-based model for non-autoregressive generation
Controlled text revision In the style transfer task, we additionally condition generation on an initial state $x^{(0)}$, and the objective is to modify $x_{0}$ to adhere to the style of $y$ while preserving the original meaning of $x^{(0)}$. To do so, we employ $E_{1}$ and $E_{2}$ as before, but introduce further experts that are functions of $x$ and $x^{(0)}$ and measure meaning preservation. We note that various options are possible for this purpose and our specific choices may not be optimal in all cases. In our experiments, we add $E_{3}$ to measure of semantic similarity using SBERT since it is not restricted to sentence-level similarity Reimers and Gurevych (2019). To ensure that $x$ makes minimal revisions to $x^{(0)}$, we additionally add $E_{4}$ defined as the Hamming distance between $x$ and $x^{(0)}$, which was also employed by Mireshghallah et al. (2022).
3.3 Inference
We frame the generation problem as finding an output $x$ which minimizes energy defined by Equation (missing) 1. Although this problem is intractable, the Metropolis-Hastings (MH) algorithm can be used to obtain an approximate sample from the desired distribution. Specifically, similar to Goyal et al. (2021), we use a pre-trained masked language model (MLM) as a token-level proposal distribution. Using a MLM enables bi-directional context to be captured by the proposal distribution. In our episodic variant of the EBM, the state of the sampler consists of more than one document. At each step, we sample one of the documents $i$ for a MH update uniformly at random, and make proposals according to the MLM conditioned on $x_{i}$, but evaluate $E_{2}(x,y)$ based on the entire sampler state. Thus, the energy function captures similarities on the entire state as opposed to a single document.
4 Experiments
4.1 Style Control
In our first experiment, we control the style of generated text using a pretrained representation of an author’s style, thereby learning to generate text in an arbitrary writing style.
4.1.1 Metrics
Fine-grained control on an author specific level is challenging to evaluate due to the subtlety of features which identify an author. In practice, authorship attribution requires trained forensic linguists to identify such distinguishing features. Therefore, our evaluation of control success relies on automatic metrics. To avoid concerns about gaming certain metrics, we include multiple automatic metrics for each text attribute that is measured. We measures the overall quality of generated text through fluency in addition to particular features (e.g. $n$-gram overlap, semantic meaning, style consistency) in generated text Celikyilmaz et al. (2020). We also consider further downstream tasks to evaluate the quality of style altered text, like author detection (§missing 4.3) and LM generated text detection (§missing 4.4).
Style similarity
To measure how well generated text matches a target style, we adapt previous work (discussed in §missing 2) as automatic evaluation tools. We consider “Universal Author Representations” (UAR) Rivera-Soto et al. (2021) and “Content Independent Style Representations” (CISR) Wegmann et al. (2022). [^6] We use two open source models for evaluation. We train our own UAR model using code provided by https://github.com/LLNL/LUAR and the CISR checkpoint is published here: https://huggingface.co/AnnaWegmann/Style-Embedding These pre-trained embeddings measure style overlap between generated and reference text samples. We report cosine similarity between reference and generated text embeddings.
Fluency
Beyond satisfying style-specific constraints, generated text should remain coherent. We report perplexity for reference and generated text under GPT-2 Radford et al. (2019). We emphasize however that fluency is, to some extent, at odds with the goal of introducing author-specific style, as evidenced by the perplexity assigned to human-written comments in Table 1. That is, if the target style samples have a high perplexity under GPT-2, it is reasonable to expect a well formed style controlled generation to also have a high perplexity. Indeed, we find that generic LMs, like GPT-3, produce very low fluencies in sharp tension with higher human reference fluencies.
Diversity
A straightforward way to mimic the writing style of a given sample $y$ is to directly copy portions of the evidence text unchanged. Useful downstream applications (e.g. writing assistants, text debiasing) are not useful if they are repetitive, and so we are interested in sampling diverse outputs which satisfy the given style constraints. We consider a Jaccard similarity over the unique 1 and 2-grams in the reference and generated text samples to measure the overlap. A lower similarity suggests a more favorable generation with less direct copying.
4.1.2 Experimental Setup
Evaluation datasets
We evaluate the effectiveness of our decoding strategy on 1000 authors contributing to four unique Reddit subreddits. We consider Reddit data collected through the Pushshift API Baumgartner et al. (2020a) and compile a test split for each of our 4 subreddits: /r/wsb, /r/AskHistorians, /r/news, and /r/australia; performance on all four results are reported together in Table 1. These subreddits are selected for their unique and distinctive styles. We use an additional split from /r/wsb to validate our methods (i.e. to select optimal decoding hyperparameters). For each author, we compile $N$ text samples as the source of style evidence. In all experiments except for Figure 1, we set $N=16$ to balance the quality of resulting style representations and cost of compiling data following previous work Andrews and Bishop (2019). Figure 1 illustrates part of this trade-off, with improved style controlled generation for larger $N$ values.
Language models
Our proposed methods operate on a frozen underlying LM; we do not perform any fine-tuning. Across all experiments we use variants of OPT and MPT-7B to generate text Zhang et al. (2022); Team (2023). [^7] A single NVIDIA V100 GPU was sufficient to store both the control model and language model, except for 7B parameter experiments which require a second V100 GPU. For EBMs, we use RoBERTa-base as our MLM for the token-level proposal distribution.
Future Regressors
We train the forward looking regressor using a single V100 GPU, a batch size of 64, and a learning rate of $1e^{-6}$ for 100k steps. In our experiments we apply this re-scoring procedure to OPT-125m to capture author specific fluencies, and we evenly weight likelihoods from OPT and the forward regressor to compute re-scored sequence level likelihoods.
$\uparrow$ Success
Model Size UAR CISR $\downarrow$ Fluency $\downarrow$ J-1 $\downarrow$ J-2
Human Reference 0.898 0.864 1356.88 0.161 0.023
Prompting (GPT-3) 175B 0.815 0.649 62.73 0.152 0.031
Prompting (GPT-3.5) 0.782 0.617 945.71 0.110 0.015
Future Regressor (OPT) 350M + 82M 0.837 0.727 373.83 0.110 0.012
1.3B + 82M 0.826 0.707 362.23 0.108 0.013
2.7B + 82M 0.836 0.730 366.88 0.102 0.009
6.7B + 82M 0.832 0.721 348.97 0.100 0.007
Future Regressor (MPT) 7B + 82M 0.829 0.715 543.84 0.101 0.009
[…]
Prompting Baseline
As a comparative state of the art baseline, we use few-shot prompting to prompt GPT-3 Brown et al. (2020) and GPT-3.5 for stylistic generations. We utilize the prompt-learning framework described by Liu et al. (2021) to generate outputs which can replicate an author’s style. We provide a template for our prompt, where each writing sample in the prompt is truncated to 32 tokens.
@@ 4.1.2 Experimental Setup
For the GPT-3 baselines, we use the largest model, Davinci, with 175 billion parameters Brown et al. (2020); for GPT-3.5 baselines, we use the gpt-3.5-turbo-0613 snapshot. We chose generation hyperparameters based on the best results on validation data from r/wsb.
With GPT-3, we use temperature of 1.0 and frequency penalty of 2; with GPT-3.5, we use temperature of 2.0 and frequency penalty of 2. We terminate generations after 32 tokens.
4.1.3 Stylistic Generation
Our style control results are shown in Table 1, where we compare our proposed non-autoregressive decoding strategy (denoted EBM), ablations consisting of just the adapted LM using future regressors, and prompting-based methods using state of the art LMs. The EBM results in the final two rows use the future regressor outputs as initializations for revision  [^8] We did experiment with revising GPT outputs as well, but found that revising the future regressor output yielded better results. This initialization is supported by our use of future regressors to balance fluency in our energy function.. The first row shows metrics for "gold" style matches, i.e. additional held out text samples written by the same human author are used for comparison. Our proposed decoding strategy performs competitively despite the fact that the baseline LMs are much larger and have undergone steps like instruction tuning in the case of GPT-3.5 Ouyang et al. (2022b). Under the UAR style metric, our proposed future regressor method outperforms baseline LMs using in-context learning. When the outputs of the future regressor are revised using the EBM text revision method described in §missing 3.2999For these experiments, we iteratively sample for 5 epochs, where an epoch iterates for the number of tokens in the longest sentence in the batch., it outperforms the prompting method on both success metrics, also shown in Table 1.
4.1.4 Style Vector Interpolation
@@
4.2 Style Transfer
In our second experiment, building on our style control experiments we explore whether we can produce text in an arbitrary writing style while preserving the meaning of the original text. To control both style and meaning, we use the EBM style-controlled revision method described in §missing 3.2.
4.2.1 Metrics
The same metrics used to evaluate style control in §missing 4.1.1 can be used for style transfer. However, since style transfer necessitates the preservation of meaning, we introduce another metric to measure this.
@@ 4.2.1 Metrics
Since our experiments involve short documents, we consider semantic search models which provide a document-wide notion of semantic similarity. Specifically we employ (1) all-mpnet-base-v2, a high-performance SBERT model and (2) GTR Ni et al. (2021), a large dual encoder trained for semantic search. We emphasize that the SBERT model used for evaluation is distinct from the all-MiniLM-L6-v2 SBERT model used in the EBM.
4.2.2 Experimental Setup
For style transfer we create a dataset pairing Reddit comments with arbitrary target styles. We randomly select three author styles from each of the 4 subreddits specified in §4.1.2 and /r/casualUK for a total of 15 target styles. Since our author styles are derived from the author’s comment history, we can pair these comments with the other author styles in a round-robin manner. We exclude pairings between comments and styles that co-occur in the same subreddit, yielding a total of $2880$ pairs. We construct a separate dataset for pairings that occur within the same subreddit, yielding $480$ total pairs.
Meaning Preservation
For EBMs, as proposed in §missing 3.2 we include a “meaning preservation” expert. For this expert, we use the all-MiniLM-L12-v2 Sentence Transformers model.
@@ 4.2.2 Experimental Setup
For in-context style transfer, we use a 2-shot variation of the approach described by Patel et al., where text is first paraphrased into a neutral style before being rewritten to match the target style. We keep the same hyperparameters as in §missing 4.1.2, except that we extend the maximum length of most generations to 64 and reduce any temperatures set at 2.0 to 1.0.
4.2.3 Style Transfer Results
@@
Model UAR CISR SBERT GTR $\downarrow$ Fluency
Random 0.729 0.334 0.042 0.438 437.14
Prompting (GPT-3) 0.811 0.752 0.650 0.784 189.97
Prompting (GPT-3.5) 0.774 0.596 0.760 0.853 87.84

### 04 Mar 2024 v2

@@ 4.2 Experimental Setup
Human Reference 0.852 0.864
GPT-3 175B 0.755 0.649 110.2%
GPT-3.5 0.742 0.560 77.23%
GPT-4 0.767 0.637 46.93%
Future Regressor (OPT) 350M + 82M 0.791 0.727 706.14%
1.3B + 82M 0.789 0.707 625.97%
2.7B + 82M 0.759 0.730 128.92%
6.7B + 82M 0.767 0.721 78.38%
Future Regressor (MPT) 7B + 82M 0.788 0.715 95.33%
MuCoLa Kumar et al. (2022) 125M + 82M 0.709 0.596 677.24%
StyleMC (Revise Future Regressor) 3B + 1.3B + 82M 0.849 0.762 322.33%
StyleMC  (Revise GPT-4) 0.838 0.736 39.87%
Table 1: Test results on all four subreddit test splits. The proposed future regressor approach outperforms both prompting approaches on the target control metric (UAR) and the secondary style metric (CISR). Both our proposed model and MuCoLa revise the output of the OPT-350M future regressor model. We additionally use our method to revise GPT-4 outputs, finding significant improvements. In both cases, revising outputs with our method achieves the highest performance, approaching the human reference. We omit model sizes for GPT-3.5 and GPT-4, as they are unknown. A paired sign test of the differences between our proposed method and GPT-4 is significant at at least the $p<10^{-7}$ level for all metrics.
Baselines
We use few-shot prompting with GPT-3 Brown et al. (2020), GPT-3.5, and GPT-4. Since stylistic generation does not impose a constraint on semantic similarity, we found that a simple prompt was sufficient, compared to the schemes proposed by Reif et al. (2021) and Patel et al. (2022) for style transfer. We provide a template for our prompt, where each writing sample in the prompt is truncated to 32 tokens.
@@
For the GPT-3 baselines, we use the largest model, Davinci, with 175 billion parameters Brown et al. (2020); for GPT-3.5 baselines, we use the gpt-3.5-turbo-0613 snapshot; for GPT-4 we use gpt-4-0613. For all models, we use a temperature of 1.0 and frequency penalty of 2, terminating generations after 32 tokens.
We also compare our proposed text revision approach to MuCoLa Kumar et al. (2022), a sampling procedure that uses gradients to optimize over differentiable constraints. We report the results in Table 1. A limitation of this approach is that all models must share the same embedding table; we use a LUAR style model trained to use the OPT-125M embedding table, which obtains within 4.7% of the performance of the style model. We construct an energy function with the goal of maximizing the cosine similarity between the target style embedding and the style embedding of a single output from our future regressor. We tune hyperparameters on the validation split of our dataset and use the best hyperparameters reported for the sentiment task in the original paper with several exceptions: we use the weighted sum selection criteria with a weight of 0.001 on the language model and 0.999 on the style model, a threshold of -5, and maximum length of 32 tokens.
4.3 Style-Controlled Generation
Table 1 compares the proposed approach to ablations consisting of just the adapted LM using future regressors, and prompting-based methods using state of the art LMs. The results in the final two rows use our author-adapted LMs to sample initializations for revision. [^10] We did experiment with revising GPT outputs as well, but found that revising the future regressor output yielded better results. This initialization is supported by our use of future regressors to balance fluency in our energy function. The first row shows metrics for “gold" style matches, i.e. additional held out text samples written by the same human author are used for comparison. Our proposed decoding strategy performs competitively despite the fact that the baseline LMs are much larger and have undergone steps like instruction tuning in the case of GPT-3.5 Ouyang et al. (2022). Under the UAR style metric, our proposed future regressor method outperforms baseline LMs using in-context learning. When the outputs of the future regressor are revised using the EBM text revision method described in §missing 3.2111111For these experiments, we iteratively sample for 5 epochs, where an epoch iterates for the number of tokens in the longest sentence in the batch., it outperforms the prompting method on both success metrics, also shown in Table 1. We additionally demonstrate that StyleMC can improve the outputs of prompted LMs; we significantly improve control results by revising the outputs of GPT-4. Across all of our methods, we find that the amount of stylistic evidence made available directly impacts decoding performance (Figure 2).
4.4 Style Vector Interpolation
@@
5 Style Transfer Experiments
In our second experiment, building on our style control experiments we explore whether we can produce text in an arbitrary writing style while preserving the meaning of the original text. To control both style and meaning, we add a semantic similarity expert in the form of an SBERT encoder as described in §missing 3.2.
5.1 Metrics
The same metrics used to evaluate style control in §missing 4.1 can be used for style transfer. However, since style transfer necessitates preservation of meaning, we introduce another metric to measure this.
@@ 5.1 Metrics
Since our experiments involve short documents, we consider semantic search models which provide a document-wide notion of semantic similarity. Specifically we employ (1) all-mpnet-base-v2, a high-performance SBERT model and (2) GTR Ni et al. (2021), a large dual encoder trained for semantic search. Note that the SBERT model used for evaluation is distinct from the SBERT model used to guide generation (all-MiniLM-L6-v2).
Edit distance
To convey how much a sequence is changed to achieve a target style, we measure a Levenshtein distance between the initial text and final outputs from each method. Intuitively, it is desirable to make parsimonious edits to achieve a desired style, such as in the anonymization experiments in §missing 6.1.
5.2 Experimental Setup
For style transfer we create a dataset pairing Reddit comments with arbitrary target styles, where each target style consists of 16 comments from the same Reddit user and Subreddit. We select three author styles at random from each of the 4 subreddits specified in §4.2 and /r/casualUK for a total of 15 target styles. Since our author styles are derived from the author’s comment history, we can pair these comments with the other author styles in a round-robin manner. We exclude pairings between comments and styles that co-occur in the same subreddit, yielding a total of $2880$ pairs. For EBMs, as proposed in §missing 3.2 we include a “meaning preservation” expert. For this expert, we use the all-MiniLM-L12-v2 Sentence Transformers model.
@@ 5.2 Experimental Setup
For in-context style transfer, we use a 2-shot variation of the approach described by Patel et al., where text is first paraphrased into a neutral style before being rewritten to match the target style. We keep the same hyperparameters as in §missing 4.2, except that we extend the maximum length of generations to 64.
5.3 Style Transfer Results
@@
Model UAR CISR SBERT GTR $\downarrow$ $\Delta$Fluency $\downarrow$ Levenshtein
Initial State 0.660 0.435 0.042 0.438 781.78%
GPT-3 0.733 0.749 0.650 0.784 711.69% 71.68
GPT-3.5 0.701 0.590 0.760 0.853 147.76% 64.52
GPT-4 0.738 0.723 0.741 0.841 742.31% 62.60
StyleMC 0.768 0.618 0.968 0.978 701.21% 14.71
StyleMC (GPT-4 Initialization) 0.791 0.802 0.728 0.832 735.98% 17.83
   (-Noun Constraints) 0.761 0.583 0.963 0.971 807.43% 23.21
   (-Style Expert) 0.662 0.432 0.991 0.993 794.84% 1.91
   (-Semantic Expert) 0.831 0.680 0.501 0.702 1013.11% 47.67
   (-Fluency Expert) 0.755 0.599 0.968 0.976 803.98% 14.79
Table 2: Test results for style transfer. The EBMs revise human written text to the style of a different author from a different subreddit. We provide metrics for the initial text, before style transfer, as a reference point. The last column represents a Levenshtein edit distance from the model outputs to the initial text. A paired sign test of the differences between our proposed method and GPT-4 is significant (at least) at the $p<10^{-7}$ level for all metrics. We bold the best score for each metric, ignoring ablations.
In Table 2, for style-transferred text produced by each method, we report the extent to which the target style is achieved (UAR, CISR), fluency as measured by GPT-2, and the extent of semantic preservation (SBERT, GTR). As a trivial baseline, we report the same metrics between pairs of unrelated text samples, which was achieved by comparing the style reference to the initial content (i.e. initial state). Our approach performs comparably to prompted large language models while requiring only a fraction of the number of parameters.
While searching for hyperparameters, we observed a trade-off between stylistic accuracy and content preservation. This observation is consistent with the notion that style and content cannot be completely disentangled. For instance, optimizing for content preservation may introduce stylistic features from the target content into generated text, and vice versa. We handle this trade-off by providing hyperparameters for tuning the relative importance between semantic preservation and style accuracy, providing a “knob” to tune for different applications. We show qualitative results in Table 3 and Appendix A, demonstrating that prompted LMs often fail to capture finer-grained changes necessary for style transfer. Our proposed models are able to selectively edit sequences to satisfy these constraints.
Target Style
"Reuse requires more effort/labour. People are lazy, especially in australia."
"What you learn at uni isnt rubbish but really a framework for your skills. Essentially you’re learning to learn."
"I’d like Tanya too but in general Australian just isn’t ready a female leader again."
Initial Content
"lol global warming is a fact? Please, we all know it’s not after they changed the goal post to ‘climate change’ instead"
"this brings me some hope for my -70% gains."
"so are you saying we aren’t using atrazine as herbicide for our crops?"
GPT-4
"The phrase ‘global warming’ was rebranded as ‘climate change’, which caused some to doubt its authenticity."
"This brings a glimmer of hope to my 70% losses."
"Are you saying we’re not using atrazine as a pesticide on our crops?"
Proposed (RoBERTa Proposals)
"Our climate warming’s not fact. Worse, others have realised it is not because we switched the goal post to ‘climate change’ instead"
"That gives me some hope for my -70% gains today."
"Why were we saying we aren’t doing atrazine as herbicide for the crops."
Proposed (T5-3B Proposals)
"To that end. If global warming is a fact? Ahhh, bloody heck, we all know it’s not after they had switched the goal post to just ‘climate change’ instead"
"The above has given me some hope for my -70% gains."
"So are you saying that we aren’t using atrazine as herbicide for our crops?"
Table 3: Successful style transfer from a r/wsb author to a r/australia author. Obvious stylistic behavior is highlighted in orange, and edits by our proposed models are in red font. GPT-4 is able to edit punctuation and capitalization, but fails to capture finer-grained features. T5-3B is able to replace multiple tokens at a time when necessary, as indicated in this table leading to better style transferred results.
6 Additional Experiments
6.1 Anonymization
Enabling author privacy is a promising application of StyleMC. Previous work has explored the preservation of privacy by altering identifying linguistic features associated with text Li et al. (2018). We measure success by the system’s ability to circumvent an author attribution system. In this setting, author attribution involves attempting to match text samples $Q$ (queries) and $T$ (targets) that were written by the same author. We consider a subset of the authors in the Reddit dataset to evaluate attribution capabilities. Our sample consists of 180 authors, and results in 32,400 binary comparisons. Given a user’s history made up of $N$ posts, we take the first $N/2$ posts and establish a query ($Q)$, and the second $N/2$ posts to establish a target ($T$). In our experiments we use $N=16$. Using our proposed style transfer approach, we alter the style of each target $T$ to produced a perturbed target $T^{\prime}$. Success is measured by the decrease in performance when matching $Q\rightarrow T^{\prime}$ compared to $Q\rightarrow T$. To evaluate our approach, we consider all possible pairs of queries and targets and seek to detect matching queries and targets before and after style transfer is applied. We extract representations using UAR for each query and target sample, and compute pairwise distances to use as scores Rivera-Soto et al. (2021). A smaller score in this case indicates a higher likelihood that the two representations are from the same author. Solving the detection problem involves setting an operating point with a given rate of false positives and false negatives, the point at which the two rates are equal is known as the equal error rate. A lower value indicates a better detection result. The results in Table 4 show that our procedure successfully reduces the detection rate through style transfer.
[…]
GPT-3 0.116 $\rightarrow$ 0.311 71.68
GPT-3.5 0.116 $\rightarrow$ 0.278 64.52
GPT-4 0.116 $\rightarrow$ 0.324 62.60
Proposed 0.116 $\rightarrow$ 0.363 14.71
Table 4: Extent of anonymization after style transfer. An increasing EER via style transfer indicates improved anonymization. The last column is a Levenshtein distance between the model output and original text. StyleMC is able to more effectively anonymize writing with fewer edits.
6.2 Detection of generated text
Strategy GPT-3 StyleMC
Zero-shot 0.676 0.547
In-domain Training 0.972 0.831
Table 5: Detection accuracy for text sampled from GPT-3 and our proposed decoding strategy. Each split consists of 250 real and 250 fake text samples.
To construct a dataset for this task, we follow a strategy used by OpenAI’s fake text detector AIT . Similar to our main experiments, we use the Pushshift API to collect real text samples from 10,000 Reddit users, ensuring that each sample has at least 16 posts Baumgartner et al. (2020). We concatenate this data to create a prompt, and allow OPT-6.7B to generate follow on fake text for the prompt. The resulting dataset consists of 10,000 human written text samples and 10,000 machine generated outputs associated with those prompts. Additionally, we construct two more datasets which include 500 GPT-3 samples and 500 samples from our proposed EBM strategy to demonstrate improved detectability when in-domain data is considered. We fine-tune a RoBERTa base model Liu et al. (2019) on these datasets for 10 epochs on a single V100 GPU using a learning rate of 2e-5 and AdamW optimizer Loshchilov and Hutter (2019).
7 Related Work
Effective text style transfer is important for many downstream applications such as writing assistants, personalized NLP systems, text simplification, detoxifying and debiasing text Jin et al. (2022). Interest in the task has led to many datasets spanning various types of styles and domains Briakou et al. (2021); Madaan et al. (2020); Rao and Tetreault (2018) and approaches Prabhumoye et al. (2018); Krishna et al. (2020); Riley et al. (2021); Hallinan et al. (2023). However, these approaches largely focus on coarse level styles (e.g. formality, politeness, simplicity) rather than fine-grained styles which may contain any combination of coarse styles. For finer-grained style transfer, Riley et al. (2021) propose a few-shot strategy using learned style vectors to autoregressively decode text. Our work differs by using a pre-existing encoder for style vectors Rivera-Soto et al. (2021) and incorporating bidirectional context during inference. In high resource settings, parallel corpora can be leveraged to directly learn relationships between styles Jhamtani et al. (2017), however having such datasets for arbitrary authors is not realistic. Additionally, recent interest in prompting large language models has facilitated style transfer from arbitrary authors using in-context learning Reif et al. (2021); Patel et al. (2022). We use similar prompting strategies as comparable baselines for our approach.
Research towards controllable text generation has focused on fine-tuning approaches, discriminator guided decoding, and more recently on large-language model prompt engineering. Fine-tuning approaches condition a language model on a given control attribute. For a control attribute $c$, the language model is trained to predict the probability of the next word $p(x\mid c)$. This probability can be directly modelled as in the case of CTRL which uses an initial control prefix to guide decoding (Keskar et al., 2019). However, CTRL requires re-training a LM any time a new control code is proposed. One way to avoid training from scratch is to approximate the probability $p(x\mid c)$ as $p(c\mid x)p(x)$. Here $p(x)$ can be modeled by a pre-trained language model and $p(c\mid x)$ can be modeled by a simple discriminator. Rather than training an entire language model, only the discriminator would need to be trained (Dathathri et al., 2019; Krause et al., 2020; Yang and Klein, 2021). However, as noted in §missing 1, control attributes perform poorly on finer-grained tasks, motivating the use of control vectors instead.
8 Conclusion
With StyleMC we have demonstrated the ability to guide the style of generated text using author representations, which capture fine-grained aspects of writing style on the basis of a small writing sample. We develop a novel sequence-level model for this purpose, consisting of an author-adapted LM and non-autoregressive inference procedure. The proposed approach outperforms large intruction-tuned LMs at guiding generated text towards the desired attributes.
Limitations
The main limitation of our study is the reliance on automatic evaluation metrics. To avoid relying on any single automatic metric, we include a diverse set of evaluation strategies, particularly the interpolation experiments in §missing 4.4 that focus on interpretable stylistic attributes. The success of the interpolation experiments provide support both the effectiveness of the style representations as well as our ability to generate text in the target style. In the case of coarse style transfer problems like formality and sentiment, non-expert human annotators can perform the task and therefore be used to complement automatic metrics. However, authorship attribution requires trained forensic linguists, an avenue which we decline to pursue in this work, both for cost reasons and to avoid setting a precedent that may detract from future work in this area. Similar to previous efforts in controllable generation, the proposed approach uses a discriminative model to guide generation, and success at control is reliant on the quality and availability of appropriate training data to estimate that model. In our experiments, we rely on representations of author style that are trained on large amounts of anonymous social media content and are highly discriminative of authorship Rivera-Soto et al. (2021). However, social media data may contain various biases, such as a prevalence of English over other languages, as well as biases owing to the sample sizes of various demographic groups relative to the population.
Broader Impact
This paper pushes the state of the art in style-controlled text generation, which enables a number of downstream applications, such as writing assistants, anonymization (e.g., for political dissidents), and personalized NLP more broadly, such as for under-represented groups. Another interesting potential application area is machine-translation, tailored to an author’s particular style. We are also excited about potential applications of style-controlled generation to data augmentation and synthetic data creation with LLMs, which may otherwise suffer from lack of diversity relative to real data composed by a variety of authors with distinct styles. At the same time, as with most technologies there is potential for abuse. In §missing 6, we address one way methods discussed here may be abused: defeating machine-text detectors. We explore a mitigation scheme involving retraining the detector on style-controlled outputs (Table 5), showing that this results in drastic improvements in detection accuracy. Few-shot detection approaches could also be effective in mitigating abuses of the proposed method Soto et al. (2024).
Acknowledgements
Appendix A Additional Qualitative Results
Table 6 demonstrates additional style transfer model outputs, here the transfer occurs in the opposite direction of Table 3. In one case, both of our proposed models fail to make any edits to the initial text during the sampling procedure. We again observe GPT-4 producing fluent and readible text, but with no noticeable style specific edits.
Target Style
"omg thats amazing!! where’d it go? do it again"
"this brings me some hope for my -70% gains"
"lol global warming is a fact? Please, we all know it’s not after they changed the goal post to ‘climate change’ instead"
Initial Content
"Reuse requires more effort/labour. People are lazy, especially in australia."
"What you learn at uni isnt rubbish but really a framework for your skills. Essentially you’re learning to learn."
"I’d like Tanya too but in general Australian just isn’t ready a female leader again."
GPT-4
"recycling takes work. aussies especially are too lazy to do it cause they want things easy and convenient."
"the knowledge you learn in college isn’t worthless, it gives you a foundation of skills. essentially, you’re getting schooled on how to school yourself."
"i like Tanya but aus ain’t ready for another sheila in charge."
Proposed (RoBERTa Proposals)
"Reuse requires more effort/labour. People are lazy, especially in australia."
"what you learn at uni isnt rubbish sonz a framework to your skills. Essentially you learned how to learn more"
"i’d like Tanya too; good fact Australian just isn’t ready a female leader again."
Proposed (T5-3B Proposals)
"Reuse requires more effort/labour. People are lazy in australia."
"What you learn at uni isnt rubbish but really a framework for your skills. Essentially you’re learning to learn."
"tf I like Tanya so much now… but I really can’t believe she’s a feminist. Australian politics is never going to have a single male leader again."
Table 6: Style transfer from a r/australia author to r/wsb author. Obvious stylistic behavior is highlighted in orange, and edits by our proposed models are in red font. GPT-4 paraphrases the outputs but does not effectively transfer style.
Appendix B Proposal Model Variations
We experiment with several different components of our proposed model, which masks a portion of the current sequence, and samples an alternative infill. We find that the ability to produce variable length proposals to be crucial to producing high quality outputs. Table 7 demonstrates the significantly better job done by T5 at matching reference fluencies, this can be qualitatively observed in Table 3 as well. We also vary the the masking procedure, finding that sampling two tokens at a time produces the best results (Table 8).
Model $\uparrow$ UAR $\uparrow$ SBERT $\downarrow$ $\Delta$Fluency
Initial State 0.674 0.032 792.73%
RoBERTa-Base 0.766 0.949 1801.01%
T5-base 0.746 0.969 791.85
T5-large 0.752 0.965 807.21%
T5-3B 0.768 0.968 701.21%
Table 7: Style transfer results for varying proposal models. RoBERTa proposes single token edits, while T5 models may propose variable edits.
Model $\uparrow$ UAR $\uparrow$ SBERT $\downarrow$ $\Delta$Fluency
Initial State 0.674 0.032 792.73%